OpenStudy (anonymous):

If BD = 4x + 8 and AE = 7x + 30, find x. (HINT: remember BD = ½ AE or AE = 2BD) HELPP!!!!! WILL FAN AND MEDAL PLZZ HELP!

OpenStudy (ikram002p):

do u have a graph or formula or something ?

OpenStudy (anonymous):

its from like a study guide that i need help wrth

OpenStudy (anonymous):

Do you have a picture of the figure?

OpenStudy (anonymous):

it doesmt show a figure

OpenStudy (anonymous):

OK let's try this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

AE = 2BD. And AE = 7x + 30 and BD = 4x + 8 , so \(7x + 30 = 2(4x +8)\). Solve for x